Given the table of values for s and t, write down a formula expressing s in term of t for these values :

Answer by Theo(13342)

suppose your table looks like this:

               x                t

               0                3
               1                8
               2                13
               3                18
               4                23


your formula expressing t in terms of s would be t = 5s + 3

when s = 1, t is equal to 5 * 1 + 3 = 8
when s = 4, t is equal to 5 * 4 + 3 = 23

this is a linear equation, so each succeeding value of t will always be 5 more than the preceding value of t when each succeeding value of s is 1 more than the preceding value of s.

you can see this from the table.

when s = 0, the value of t will be 0 * 5 + 3 = 3.

the graph of this equation will look like this, after replacing t with y and s with x.

$$$


as you can see, this is the graph of a straight line.

the vertical dashed lines are only there to show you the value of y for each value of x.

the value of y on the graph represents the value of t.

the value of x on the graph represents the value of x.

the coordinates are in the form of (x,y).

x represents the distance the line travels horizontally in a direction parallel to the x-axis.

y represents the distance the line travels vertically in a direction parallel to the y-axis.

for example, if you look at the point (4,23) and the point (3,18), you will see that the point (4,23) is 1 unit to the right and 5 units up from the point (3,18).

the difference in the x-coordinates is 4 - 3 = 1
the difference in the y-coordinates is 23 - 18 = 5

the value of x went up 1 and the value of y went up 5.

this is all according to the equation y = 5x + 3.

when x was 3, the value of y was 18.
then x was 4, the value of y was 23.

the value of x went up 1.
the value of y went up 5.

on the graph, the value of x going up 1 unit means going to the right 1 unit.
on the graph, the value of y going up 5 units means going up 5 units.